OpenStudy (anonymous):

need help with this calc problem using substitution. 3 sqrt 2-x over x exp7. thks

OpenStudy (anonymous):

use the equation toolbox and write out the equation. What you described is unclear.

OpenStudy (anonymous):

ok \[ \int\limits_{1}^{2} \sqrt[3]{2-6} / x^{7}\]

OpenStudy (anonymous):

this is integration?

OpenStudy (anonymous):

sorry.. forgot the dx

OpenStudy (anonymous):

okay. so take out the cuberoot(2-6) outside the integral. you will have \[\sqrt[3]{2-6} \int\limits_{1}^{2}x ^{-7}dx\]. That is fairly straightforward to solve